# Comprehensive Evaluation of OCT-based Automated Segmentation of Retinal Layer, Fluid and Hyper-Reflective Foci: Impact on Diabetic Retinopathy Severity Assessment

### v3

@@ Abstract
Background: Diabetic retinopathy (DR) is a major cause of vision loss, and early detection is essential to prevent irreversible blindness. Spectral Domain Optical Coherence Tomography (SD-OCT) enables high-resolution retinal imaging, while AI-driven segmentation improves diagnostic precision. However, segmentation performance varies across models, especially for DR cases with differing severity and complex fluid and hyperreflective foci (HRF) patterns. The clinical deployment of these models remains underexplored. This study develops an active-learning-based deep learning pipeline for automated segmentation of retinal layers, fluid, and HRF, comparing state-of-the-art (SOTA) models and evaluating their impact on DR assessment.
@@
Results: SwinUNETR achieved the highest overall accuracy (DSC = 0.7719; NSD = 0.8149), while VM-UNet outperformed in specific layers. PDR showed increased OPL, fluid, and HRF thickness, whereas NPDR exhibited thickening in ONL+IS. In NPDR, thickening in RNFL, OPL, ONL+IS, and RPE correlated with reduced vision. In PDR, OS and EZ thickening and INL thinning were associated with visual impairment.
@@ I Introduction
Diabetic retinopathy (DR) is a prevalent microvascular complication of diabetes mellitus (DM) and a leading cause of vision impairment worldwide (Klein, 2007). It was estimated that approximately 20% of diabetic individuals over the age of 50 will develop DR, which, if left untreated, can progress to severe visual impairment or blindness (Teo et al., 2021). The progression of DR was influenced by several risk factors, including prolonged diabetes duration, elevated glycated hemoglobin (HbA1c) levels, hypertension, hyperlipidemia, obesity, and smoking (Klein et al., 2014)(Zhou et al., 2025)(Mori et al., 2024)(Cheung et al., 2016)(Zhang et al., 2024)(Cao et al., 2017). Clinically, DR is categorized into two primary stages: Non-Proliferative Diabetic Retinopathy (NPDR) and Proliferative Diabetic Retinopathy (PDR). NPDR represents an early stage, often asymptomatic, characterized by microvascular abnormalities that progressively compromise retinal capillary integrity. Without timely medical intervention, NPDR can advance to PDR, a more severe stage marked by pathological neovascularization due to chronic retinal ischemia. This progression increases the risk of severe complications such as vitreous hemorrhage and retinal detachment, ultimately threatening vision.
@@
Spectral Domain Optical Coherence Tomography (SD-OCT) is a cutting-edge, non-invasive imaging technique that provides high-resolution, cross-sectional visualization of retinal structures. Its real-time image acquisition capability makes it an invaluable tool for DR screening and early diagnosis (Gabriele E. Lang, 2007). SD-OCT enables the detection of subclinical retinal changes by quantifying variations in retinal thickness and identifying fluid accumulation. Accurate and reliable segmentation of retinal layers and pathological features is crucial for DR diagnosis and treatment planning.
@@
Numerous studies have explored automated segmentation techniques for retinal layer analysis. Herzog et al. proposed an edge maximization and smoothness-constrained thresholding approach to delineate retinal boundaries (Herzog et al., 2004). Chiu et al. utilized a graphical cut algorithm to minimize the weighted sum of edge paths along connected nodes, effectively segmenting retinal layers (Chiu et al., 2010). Wang et al. introduced a multi-step approach that includes artifact removal, contrast enhancement, and segmentation via level set methods, k-means clustering, and Markov random fields (MRFs) (Wang et al., 2015). Traditional machine-learning techniques have also been employed for fluid segmentation. González et al. identified dark fluid regions in OCT scans using support vector machines (SVM) and random forest classifiers (Gonzalez et al., 2013). Chen et al. applied a graph-cut classifier followed by a region-growing algorithm for cystoid macular edema (CME) segmentation (Xinjian Chen et al., 2012). However, these conventional approaches are limited by their reliance on handcrafted features and their susceptibility to performance degradation in severely diseased cases.
@@
Deep learning has emerged as a powerful alternative for automated retinal segmentation, offering greater robustness against variations in image quality and pathological abnormalities. Liu et al. utilized a ResNet-based convolutional neural network (CNN) combined with a random forest classifier for patch-wise layer segmentation (Liu et al., 2019). Kugelman et al. proposed a recurrent neural network (RNN) with a graph search framework to segment retinal layers in both healthy individuals and patients with age-related macular degeneration (AMD) (Kugelman et al., 2018). Hu et al. developed a multi-scale CNN capable of capturing different feature levels for improved segmentation accuracy (Hu et al., 2019). U-Net and its derivatives have become widely adopted among deep-learning models for medical image segmentation. U-Net’s encoder-decoder architecture, enhanced by skip connections, enables efficient spatial information preservation and mitigates vanishing gradient issues (Ronneberger et al., 2015). It has been successfully applied to retinal layer segmentation, fluid detection, and HRF analysis, achieving state-of-the-art (SOTA) performance (Roy et al., 2017, Ma et al., 2021, Tennakoon et al., 2018, Schlegl et al., 2018). Generative adversarial networks (GANs) were also used for retinal boundary augmentation and segmentation adaptation cross multiple OCT domains(Kugelman et al., 2023, Chen et al., 2023). Vision Transformers (ViTs) have recently outperformed CNNs in large-scale datasets. Unlike CNNs, which rely on local receptive fields, ViTs employ self-attention mechanisms to capture global dependencies, which is particularly beneficial for detecting diffuse fluid regions. Xue et al. implemented a Swin-Transformer-based architecture for fluid segmentation in diabetic macular edema (DME) and AMD, demonstrating superior performance over traditional CNN-based models (Xue and Du, 2024). Kulyabin et al. leveraged the Segment Anything Model (SAM) for retinal fluid segmentation, incorporating point and bounding box prompts to outperform U-Net in macular hole and fluid segmentation tasks (Kulyabin et al., 2024). Despite these advancements, most existing studies focus on either the retinal layer or fluid segmentation, with varying levels of segmentation performance on pathological clinical features. However, limited efforts are dedicated to investigating the effect of automated segmentation performance on NPDR/PDR classification or prognosis, which is crucial to evaluating their clinical translation.
@@
Studies have examined the relationship between retinal layer thickness, fluid accumulation, and DR severity. Browning et al. analyzed macular thickness across different DR severity levels and observed a correlation between macular thickening and increased risk of subclinical edema (Browning et al., 2008). Kim et al. investigated choroidal thickness alterations in DR and DME patients, reporting a significant increase in choroidal thickness as DR severity progressed from mild/moderate NPDR to PDR (Kim et al., 2013). Cho et al. assessed macular and peripapillary retinal thickness in DR subjects, identifying statistically significant differences in retinal thickness across seven anatomical regions between DR and control groups (Cho et al., 2010). Santos et al. demonstrated that fluid accumulation within the outer segment (OS) layer is significantly associated with central retinal thickness and visual impairment in DME patients (Santos T et al., 2024). These findings suggest that retinal layer thickness and fluid distribution are both reliable biomarkers for DR diagnosis and progression monitoring. However, limited efforts are dedicated to investigating the effect of automated segmentation performance on DR/PDR classification or prognosis, which is crucial to evaluating their clinical translation.
@@ II-A Data Acquisition
116 SD-OCT volumes were acquired from Sankara Nethralaya Eye Care Hospital in India. The imaging data was obtained using the Cirrus HD-OCT 5000 (Carl Zeiss Meditec, Dublin, CA, USA). Seventeen OCT volumes were captured in Macular Cube mode with a 512 × 128 pixels resolution, with the remaining 99 volumes scanned in OCTA mode at 350 × 350 pixels. Both modes covered a 6 × 6 mm² macular region centered on the fovea. Despite differences in scanning speed and resolution, Wong et al. reported no significant variation in macular thickness measurements between the two modes (Wong et al., 2024). Table I presents the demographic details of the subjects in two DR severity groups, showing no significant differences in age, diabetes duration, or visual acuity (p $>$ 0.05). However, the gender distribution differs due to the limited number of female patients in the PDR group. The variance inflation factor (VIF) is calculated for DR groups, age, gender, duration of diabetes, and visual acuity. No significant multicollinearity is found as all values are close to 1.
@@
TABLE I: Demographic information of experimental DR groups. Numerical values are presented as mean ± standard deviation (SD). The visual acuity is expressed in the logarithm of the Minimum Angle of Resolution (LogMAR). The p-values are calculated using the Welch’s t-test. The variance inflation factor (VIF) is calculated for each variable including the DR category.
@@ II-B Pre-processing
  • Motion artifacts among adjacent B-scan were corrected in both the axial and lateral directions. Axial translations were determined through cross-registration using the moving average of the central B-scan as a reference. Lateral corrections were achieved by performing registration based on the adjacent B-scans’ discrete Fourier transform (DFT). Rotational adjustments were computed by transforming translations into polar space, using the moving average of the central B-scan as a reference.
@@ II-C Active-Learning-Based Ground Truth Segmentation Annotation
Figure 1 illustrates that the active-learning-based semi-automatic segmentation follows a structured human-in-the-loop (HITL) interactive labelling workflow. SwinUNETR was used as the backbone network architecture for the active-learning workflow. The choice of architecture will only affect the efficiency of the manual labelling process, but it will not result in discrepancies in quantitative evaluations. Initially, five volumes were manually annotated from scratch. Manual segmentation was performed on every fifth B-scan, while the intermediate B-scans were interpolated under the assumption that adjacent B-scans share structural similarities. However, B-scans that exhibit significant structural changes were individually labelled and corrected. These five manually labelled volumes served as the first iteration of the training dataset to train a deep neural network (DNN) with a data split ratio of 3:1:1 for training, validation, and testing. The initially trained model was then used to generate segmentation predictions for an additional 20 volumes, which were subsequently reviewed and manually corrected. This iterative process continued with a 7:2:1 data split ratio in the subsequent iterations for training, validation, and testing, with the network being retrained on an expanded dataset each time, ensuring that all volumes undergo accurate segmentation and manual verification. Volume splits were stratified to ensure pathological cases with all label types (i.e. retinal layer, fluid, and HRF) presented in training, validation, and testing sets,
@@ II-D Automatic Segmentation Networks Architecture
  • U-Net: The most widely-used well-established medical image segmentation model that employs a CNN-based encoder-decoder architecture with skip connections (Ronneberger et al., 2015). While effective, it may struggle with high-resolution inputs due to a lack of global contextual awareness. The U-Net model is configured with a depth of five channels and three residual units.
  • SegFormer: A transformer-based architecture proposed by Xie et al., which utilizes a hierarchical transformer encoder combined with a lightweight MLP decoder to enhance feature extraction (Xie et al., 2021). The 2D variant of SwinUNETR is used for training B-scans.
  • SwinUNETR: A CNN-Transformer-composited architecture proposed by Hatamizadeh et al., which replaces the CNN-based encoder in U-Net with a Swin Transformer encoder, enabling multi-scale feature extraction through a shifted windowing mechanism. (Hatamizadeh et al., 2022). This is also the architecture that is used for the semi-automatic generation of the ground truth segmentation labels through the HITL active-learning process.
  • VM-UNet: A recently proposed novel architecture developed by Ruan et al., this model introduces a state-space model (SSM) and an asymmetric encoder-decoder structure. It models the visual data as an evolving state, efficiently capturing both local and global dynamics with a structure inspired by continuous dynamic systems, balancing computational efficiency while maintaining a global contextual view (Ruan et al., 2024).
@@ II-E Segmentation Model Training
We employed 5-fold cross-validation, stratified by DR diagnosis, with a 4:1:1 ratio for training, validation, and testing. Each input consisted of a 3-channel image constructed by three repetitions of a single B-scan. Each training B-scan was resized to $512\times 512$. To mitigate class imbalance among segmentation labels, excess Vitreous and Choroid regions were cropped. Various augmentation techniques were applied, including lateral flipping, Gaussian noise addition, contrast enhancement, rotation within the B-scan plane, and random intensity shifting.
For loss functions, we used the combinations of Dice loss, cross-entropy (CE) loss, and the L1 loss of texture differences. Given the ground truth label $y$ and predicted label $\hat{y}$, for every pixel $i$, the Dice loss is calculated as:
@@
$L_{CE}(y,\hat{y})=-\frac{1}{N}\sum_{i}\left[y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})\right]$ (2)
@@
We empirically assigned different class weights to CE loss to emphasize the class imbalance issue. Specifically, we assigned 0.1 to Vitreous and Choroid, 0.5 to the rest of the layers, and 1 to fluid and HRF. We used AdamW optimizer with CosineAnnealing scheduler with the warm restart. We adopted the distributed parallel learning supported by the PyTorch Lightning module [^1] https://lightning.ai/, with a batch size of 8 and a learning rate of 1e-4. The training was deployed on NVIDIA V100 Volta GPU allocated by Cedar Compute Canada [^2] More information can be found at: https://docs.alliancecan.ca/wiki/Cedar.
II-F Semgmentation Performance Evaluation

### v4

# Comprehensive Evaluation of OCT-based Automated Segmentation of Retinal Layer, Fluid and Hyper-Reflective Foci: Impact on Clinical Assessment of Diabetic Retinopathy Severity

@@ Abstract
Background: Diabetic retinopathy (DR) is a major cause of vision loss, and early detection is essential to prevent irreversible blindness. Spectral Domain Optical Coherence Tomography (SD-OCT) enables high-resolution retinal imaging, while AI-driven segmentation improves diagnostic precision. However, segmentation performance varies across models, especially for DR cases with differing severity and complex fluid and hyperreflective foci (HRF) patterns. The clinical deployment of these models remains underexplored. This study develops an active-learning-based deep learning pipeline for the automated segmentation of retinal layers, fluid, and HRF, comparing state-of-the-art (SOTA) models and evaluating their impact on DR assessment.
@@
Results: SwinUNETR achieved the highest overall accuracy (DSC = 0.7719; NSD = 0.8149), while VM-UNet outperformed in specific layers. PDR showed increased RNFL thickness and fluid accumulation, whereas NPDR exhibited thickening in GCL+IPL and OS. In NPDR, thickening in RNFL, INL, OPL, EZ, and the accumulation of fluid and HRF correlated with reduced vision. In PDR, RNFL thickening, GCL+IPL and ONL+IS thinning and fluid accumulation were associated with visual impairment.
@@ I Introduction
Diabetic retinopathy (DR) is a prevalent microvascular complication of diabetes mellitus (DM) and a leading cause of vision impairment worldwide [1]. It was estimated that approximately 20% of diabetic individuals over the age of 50 will develop DR, which, if left untreated, can progress to severe visual impairment or blindness [2]. The progression of DR was influenced by several risk factors, including prolonged diabetes duration, elevated glycated hemoglobin (HbA1c) levels, hypertension, hyperlipidemia, obesity, and smoking [3, 4, 5, 6, 7]. Clinically, DR is categorized into two primary stages: Non-Proliferative Diabetic Retinopathy (NPDR) and Proliferative Diabetic Retinopathy (PDR). NPDR represents an early stage, often asymptomatic, characterized by microvascular abnormalities that progressively compromise retinal capillary integrity. Without timely medical intervention, NPDR can advance to PDR, a more severe stage marked by pathological neovascularization due to chronic retinal ischemia. This progression increases the risk of severe complications such as vitreous hemorrhage and retinal detachment, ultimately threatening vision.
@@
Spectral Domain Optical Coherence Tomography (SD-OCT) is a cutting-edge, non-invasive imaging technique that provides high-resolution, cross-sectional visualization of retinal structures. Its real-time image acquisition capability makes it an invaluable tool for DR screening and early diagnosis [8]. SD-OCT enables the detection of subclinical retinal changes by quantifying variations in retinal thickness and identifying fluid accumulation. Accurate and reliable segmentation of retinal layers and pathological features is crucial for DR diagnosis and treatment planning.
@@
Numerous studies have explored automated segmentation techniques for retinal layer analysis. Herzog et al. proposed an edge maximization and smoothness-constrained thresholding approach to delineate retinal boundaries [9]. Chiu et al. utilized a graphical cut algorithm to minimize the weighted sum of edge paths along connected nodes, effectively segmenting retinal layers [10]. Wang et al. introduced a multi-step approach that includes artifact removal, contrast enhancement, and segmentation via level set methods, k-means clustering, and Markov random fields (MRFs) [11]. Traditional machine-learning techniques have also been employed for fluid segmentation. González et al. identified dark fluid regions in OCT scans using support vector machines (SVM) and random forest classifiers [12]. Chen et al. applied a graph-cut classifier followed by a region-growing algorithm for cystoid macular edema (CME) segmentation [13]. However, these conventional approaches are limited by their reliance on handcrafted features and their susceptibility to performance degradation in severely diseased cases.
@@
Deep learning has emerged as a powerful alternative for automated retinal segmentation, offering greater robustness against variations in image quality and pathological abnormalities. Liu et al. utilized a ResNet-based convolutional neural network (CNN) combined with a random forest classifier for patch-wise layer segmentation [14]. Kugelman et al. proposed a recurrent neural network (RNN) with a graph search framework to segment retinal layers in both healthy individuals and patients with age-related macular degeneration (AMD) [15]. Hu et al. developed a multi-scale CNN capable of capturing different feature levels for improved segmentation accuracy [16]. U-Net and its derivatives have become widely adopted among deep-learning models for medical image segmentation. U-Net’s encoder-decoder architecture, enhanced by skip connections, enables efficient spatial information preservation and mitigates vanishing gradient issues [17]. It has been successfully applied to retinal layer segmentation, fluid detection, and HRF analysis, achieving state-of-the-art (SOTA) performance [18, 19, 20, 21]. Generative adversarial networks (GANs) were also used for retinal boundary augmentation and segmentation adaptation cross multiple OCT domains[22, 23]. Vision Transformers (ViTs) have recently outperformed CNNs in large-scale datasets. Unlike CNNs, which rely on local receptive fields, ViTs employ self-attention mechanisms to capture global dependencies, which is particularly beneficial for detecting diffuse fluid regions. Xue et al. implemented a Swin-Transformer-based architecture for fluid segmentation in diabetic macular edema (DME) and AMD, demonstrating superior performance over traditional CNN-based models [24]. Kulyabin et al. leveraged the Segment Anything Model (SAM) for retinal fluid segmentation, incorporating point and bounding box prompts to outperform U-Net in macular hole and fluid segmentation tasks [25]. Despite these advancements, most existing studies focus on either the retinal layer or fluid segmentation, with varying levels of segmentation performance on pathological clinical features. However, limited efforts are dedicated to investigating the effect of automated segmentation performance on NPDR/PDR classification or prognosis, which is crucial to evaluating their clinical translation.
@@
Studies have examined the relationship between retinal layer thickness, fluid accumulation, and DR severity. Browning et al. analyzed macular thickness across different DR severity levels and observed a correlation between macular thickening and increased risk of subclinical edema [26]. Kim et al. investigated choroidal thickness alterations in DR and DME patients, reporting a significant increase in choroidal thickness as DR severity progressed from mild/moderate NPDR to PDR [27]. Cho et al. assessed macular and peripapillary retinal thickness in DR subjects, identifying statistically significant differences in retinal thickness across seven anatomical regions between DR and control groups [28]. Santos et al. demonstrated that fluid accumulation within the outer segment (OS) layer is significantly associated with central retinal thickness and visual impairment in DME patients [29]. These findings suggest that retinal layer thickness and fluid distribution are both reliable biomarkers for DR diagnosis and progression monitoring. However, limited efforts are dedicated to investigating the effect of automated segmentation performance on DR/PDR classification or prognosis, which is crucial to evaluating their clinical translation.
@@ II-A Data Acquisition
116 SD-OCT volumes were acquired from Sankara Nethralaya Eye Care Hospital in India. The imaging data were obtained using the Cirrus HD-OCT 5000 (Carl Zeiss Meditec, Dublin, CA, USA). Seventeen OCT volumes were captured in Macular Cube mode with a 512 × 128 pixel resolution, while the remaining 99 volumes were scanned in OCTA mode at 350 × 350 pixels. Both modes covered a 6 × 6 mm² macular region centered on the fovea. Despite differences in scanning speed and resolution, Wong et al. reported no significant variation in macular thickness measurements between the two modes [30]. Table I presents the demographic details of the subjects in two DR severity groups, showing no significant differences in age, diabetes duration, or visual acuity (p $>$ 0.05). However, the gender distribution differs due to the limited number of female patients in the PDR group. The variance inflation factor (VIF) is calculated for DR groups, age, gender, duration of diabetes, and visual acuity. No significant multicollinearity is found as all values are close to 1.
@@
TABLE I: Demographic information of experimental DR groups. Numerical values are presented as mean ± standard deviation (SD). The visual acuity is expressed in the logarithm of the Minimum Angle of Resolution (LogMAR). The p-values are calculated using the Welch’s t-test. The variance inflation factor (VIF) is calculated for each variable, including the DR category.
@@ II-B Pre-processing
  • Motion artifacts among adjacent B-scans were corrected in both the axial and lateral directions. Axial translations were determined through cross-registration using the moving average of the central B-scan as a reference. Lateral corrections were achieved by performing registration based on the adjacent B-scans’ discrete Fourier transform (DFT). Rotational adjustments were computed by transforming translations into polar space, using the moving average of the central B-scan as a reference.
@@ II-C Active-Learning-Based Ground Truth Segmentation Annotation
Figure 1 illustrates that the active-learning-based semi-automatic segmentation follows a structured human-in-the-loop (HITL) interactive labelling workflow. SwinUNETR was used solely to generate initial auto-segmentations, which were subsequently corrected by human annotators, thereby improving labelling efficiency. However, all models, including SwinUNETR, were trained and evaluated from scratch using the finalized, expert-corrected labels. As such, the initial use of SwinUNETR does not introduce any bias or discrepancies in the quantitative evaluation of model performance. Initially, five volumes were manually annotated from scratch. Manual segmentation was performed on every fifth B-scan, while the intermediate B-scans were interpolated under the assumption that adjacent B-scans share structural similarities. However, B-scans that exhibited significant structural changes were individually labelled and corrected. These five manually labelled volumes served as the first iteration of the training dataset to train a deep neural network (DNN) with a data split ratio of 3:1:1 for training, validation, and testing. The initially trained model was then used to generate segmentation predictions for an additional 20 volumes, which were subsequently reviewed and manually corrected. This iterative process continued with a 7:2:1 data split ratio in subsequent iterations for training, validation, and testing, with the network being retrained on an expanded dataset each time. This ensured that all volumes underwent accurate segmentation and manual verification. Volume splits were stratified to ensure pathological cases with all label types (i.e. retinal layer, fluid, and HRF) were presented in training, validation, and testing sets.
@@ II-D Automatic Segmentation Networks Architecture
  • U-Net: The most widely-used, well-established medical image segmentation model that employs a CNN-based encoder-decoder architecture with skip connections [17]. While effective, it may struggle with high-resolution inputs due to a lack of global contextual awareness. The U-Net model is configured with a depth of five channels and three residual units.
  • SegFormer: A transformer-based architecture proposed by Xie et al., which utilizes a hierarchical transformer encoder combined with a lightweight MLP decoder to enhance feature extraction [31]. The 2D variant of SwinUNETR is used for training B-scans.
  • SwinUNETR: A CNN-Transformer-composited architecture proposed by Hatamizadeh et al., which replaces the CNN-based encoder in U-Net with a Swin Transformer encoder, enabling multi-scale feature extraction through a shifted windowing mechanism. [32]. This is also the architecture that is used for the semi-automatic generation of the ground truth segmentation labels through the HITL active-learning process.
  • VM-UNet: A recently proposed novel architecture developed by Ruan et al., this model introduces a state-space model (SSM) and an asymmetric encoder-decoder structure. It models the visual data as an evolving state, efficiently capturing both local and global dynamics with a structure inspired by continuous dynamic systems, balancing computational efficiency while maintaining a global contextual view [33].
@@ II-E Segmentation Model Training
We employed 5-fold cross-validation, stratified by DR diagnosis, with a 4:1:1 ratio for training, validation, and testing. Each input consisted of a 3-channel image constructed by three repetitions of a single B-scan. Each training B-scan was resized to $512\times 512$. To mitigate class imbalance among segmentation labels, the excess Vitreous and Choroid regions were cropped. Various augmentation techniques were applied, including lateral flipping, Gaussian noise injection($\mu=0,\,\sigma^{2}=0.01$), random contrast enhancement ($\gamma\in(0.5,4.5)$), rotation within the B-scan plane (± 20°), random zooming ($(0.5,1.5)$) and random intensity shifting (± 10).
For loss functions, we used combinations of Dice loss, cross-entropy (CE) loss, and L1 loss for texture differences. Given the ground truth label $y$ and predicted label $\hat{y}$, for every pixel $i$, the Dice loss is calculated as:
@@
$L_{CE}(y,\hat{y})=-\frac{1}{N}\sum_{i}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$
@@
We empirically assigned different class weights to CE loss to emphasize the class imbalance issue. Specifically, we assigned 0.1 to Vitreous and Choroid, 0.5 to the rest of the layers, and 1 to fluid and HRF. We used AdamW optimizer with CosineAnnealing scheduler with the warm restart. We adopted the distributed parallel learning supported by the PyTorch Lightning module [^1] https://lightning.ai/, with a batch size of 8 and a learning rate of 1e-4. The training was deployed on an NVIDIA V100 Volta GPU allocated by Cedar Compute Canada [^2] More information can be found at: https://docs.alliancecan.ca/wiki/Cedar.
II-F Segmentation Performance Evaluation